OpenStudy (anonymous):

y=2x+5 y=-2x-11 solve each system by substitution

OpenStudy (anonymous):

2x+5=-2x-11 4x=-16 x=-4 y=2(-4)+5=-3 (-4.-3)

OpenStudy (anonymous):

thankss